OpenStudy (anonymous):

Given the function g(x) = 8x − 2, find g(−2)? Options are 14,0.5,-18,and -14. I thought it was C or D but just want to be sure.

OpenStudy (anonymous):

since g(x)=8(x)-2 then g(-2)=8(-2) - 2 = ????

OpenStudy (anonymous):

do the calculation

OpenStudy (hba):

none of the options the answer is -18